OpenStudy (anonymous):

Will give out medals. Edith spends 8 hours a day working in her office and 1 hour in the office taking lunch. She wrote this equation to represent the total time she spends in the office in one day. t = 8 + 1 Edith works 5 days a week. Which equation would represent the total time Edith spends in the office in one week?

OpenStudy (anonymous):

A. t = (5 • 8) + 1 B. t = 5 • (8 + 1) C. t = 8 + (5 • 1) D. t = 5 + (8 + 1)

OpenStudy (mathmale):

I'd suggest you differentiate between t (the total time spent per day) and T (the total time spent per week. How many days per week do you expect that this Edith works? if t=8+1, then what is an appropriate expression for T?

OpenStudy (anonymous):

c

OpenStudy (mathmale):

Please explain your reasoning. Then step back and ask yourself whether your reasoning in choosing an answer makes sense.

OpenStudy (mathmale):

How many days per week do you expect that this Edith works? I asked you this earliler.